OpenStudy (anonymous):

Find the equation of the tangent line to the curve (a lemniscate) 2(x^2+y^2)2=25(x^2−y^2) at the point (−3,1). The equation of this tangent line can be written in the form y=mx+b where m is:

OpenStudy (tkhunny):

There's an implicit derivative in your future.

OpenStudy (anonymous):

yes, I know that part. do i need to simp;ify first then F'?

OpenStudy (tkhunny):

Why? If you solve first, it won't be implicit. 2*2(x^2 + y^2)(2x + 2yy') = 2x - 2yy' Now solve for y'.

OpenStudy (anonymous):

how did you get this?

OpenStudy (anonymous):

where did 25 go?

OpenStudy (tkhunny):

I lost it. Please put it back. Sorry about that. End of a line confusion.

OpenStudy (anonymous):

\[f(x)=2(x^{2}+y^{2})^2=25(x^{2}-y^{2})\]

OpenStudy (anonymous):

This is the equation :9

OpenStudy (tkhunny):

No. it's not. You invented the f(x). The implicit derivative is exactly as I stated it, excepting the return of the 25. 2*2(x^2 + y^2)(2x + 2yy') = 25(2x - 2yy') Solve for y'.

OpenStudy (anonymous):

why its 2*2 in the front?

OpenStudy (anonymous):

Im confused. how did you get your equation?

OpenStudy (tkhunny):

It's an implicit dervative. Assume y = f(x) and use lots of chain rule.

OpenStudy (tkhunny):

Have you done an implicit derivative?

OpenStudy (anonymous):

prudct rule?

OpenStudy (anonymous):

No, i am not sure.

OpenStudy (tkhunny):

(d/dx)(xy) = xy' + y -- Product and Chain Rule (d/dx)(y^2) = 2yy' = Power and Chain Rule Keep in mind that we are assuming y = f(x).

OpenStudy (anonymous):

Yes! thanks for the fact. this is what exactly I do not understand. product and chain rule. how do i find out? I know the rule but when it comes to double or mixed of two rules I get confused. I have test next week, please help me

OpenStudy (tkhunny):

Well, if this is your first exposure to the implicit derivative, we are in trouble. I have no confidence that we can cover the required territory in this forum.